# UniRGB-IR: A Unified Framework for RGB-Infrared Semantic Tasks via Adapter Tuning

### v2

@@ Abstract
Semantic analysis on visible (RGB) and infrared (IR) images has gained attention for its ability to be more accurate and robust under low-illumination and complex weather conditions. Due to the lack of pre-trained foundation models on the large-scale infrared image datasets, existing methods prefer to design task-specific frameworks and directly fine-tune them with pre-trained foundation models on their RGB-IR semantic relevance datasets, which results in poor scalability and limited generalization. In this work, we propose a general and efficient framework called UniRGB-IR to unify RGB-IR semantic tasks, in which a novel adapter is developed to efficiently introduce richer RGB-IR features into the pre-trained RGB-based foundation model. Specifically, our framework consists of a RGB-based foundation model, a Multi-modal Feature Pool (MFP) module and a Supplementary Feature Injector (SFI) module. The MFP and SFI modules cooperate with each other as an adapter to effectively complement the RGB-based features with the rich RGB-IR features. During training process, we freeze the entire foundation model to inherit prior knowledge and only optimize the proposed adapter. Furthermore, to verify the effectiveness of our framework, we utilize the vanilla vision transformer (ViT-Base) as the pre-trained foundation model to perform extensive experiments. Experimental results on various RGB-IR downstream tasks demonstrate that our method can achieve state-of-the-art performance. The source code and results are available at https://github.com/PoTsui99/UniRGB-IR.git
Index Terms:
RGB-IR Semantic Tasks, Multi-modal Fusion, Adapter, Vision Transformer.
I Introduction
@@
Image semantic analysis in the visible images is a common practice in computer vision and has been widely used in a variety of vision tasks [1, 2, 3]. However, visible cameras have proved to be difficult in providing reliable imaging [4, 5], due to their limited bandwidth across the spectrum. Particularly under low-illumination conditions and adverse weather scenarios, RGB images suffer from limited informative content. Therefore, infrared (IR) modality with stronger low-light adaptability is used as additional information to supplement RGB modality [4, 6]. Afterwards, the joint use of RGB and IR images has been applied in more and more semantic analysis tasks.
@@
With the rapid advancements in physical computing capabilities, more and more general-purpose foundation backbones [7, 8] pre-trained on large-scale RGB-based datasets (e.g., ImageNet [9] and COCO [10]) have been designed for various tasks. Since these pre-trained base models implicitly encode a lot of prior knowledge, it is believed that pre-trained models can promote the improvement of downstream task performance and speed up training convergence. Therefore, fine-tuning the pre-trained foundation models [11, 12] on their semantic relevance datasets has become a common paradigm to improve the performance on specific downstream tasks.
Figure 1: Existing full fine-tuning methods vs. our UniRGB-IR framework. (a) Existing methods use pre-trained RGB-based foundation models and fully fine-tune them on their RGB-IR semantic relevance datasets. (b) We utilize the Adpater [13] to propose a new unified framework called UniRGB-IR, which can efficiently introduce richer RGB-IR features into the pre-trained foundation model for downstream tasks.
@@
However, due to the lack of pre-trained foundation models on the large-scale infrared image datasets, for RGB-Infrared semantic tasks, a straightforward and prevailing way [14, 15] is to use pre-trained RGB-based models and fine-tune them on their RGB-IR semantic relevance datasets as shown in Figure 1 (a). For example, RGB-IR object detectors [14, 16] utilize RGB-based models as strong baselines and fine-tune them to extract RGB and IR features. SwinNet [17] is explored based on the RGB-based transformer to extract hierarchical features of each modality for the detection of salient RGB-IR objects. Similarly, Ha et al. [18] establish a novel baseline by incorporating infrared modality features into an RGB-based framework and fine-tune it on the RGB-IR semantic segmentation benchmark. Among the efforts to achieve competitive results, these approaches still suffer from two major issues:
(1) Different semantic tasks require customized model structures, resulting in poor model versatility and excessive parameter storage.
(2) Full fine-tuning strategy impairs the prior knowledge encoded in the foundation model, which limits the generalization potential of the fine-tuned foundation model.
@@
Drawing inspiration from recent advances in adapters [13, 19], which are initially used to yield an extensible model to effectively exploit the potential representation of foundation models in the natural language processing (NLP) field, we develop an adapter to dynamically introduce richer RGB-IR features into the pre-trained RGB-based foundation model. In our adapter, a feature extractor is designed to obtain rich RGB-IR features and a feature injector is proposed to adaptively introduce the required features for the foundation model. Without modifying the original RGB-based foundation model, the powerful pre-trained weights can be directly frozen to accelerate training convergence and achieve efficient fine-tuning on downstream tasks. Therefore, in this paper, we propose an efficient and universal foundation model for RGB-IR semantic tasks. In particular, we are the first to construct a Unified framework for RGB-IR semantic tasks, called UniRGB-IR, as shown in Figure 1 (b).
Specifically, we design two core modules in our UniRGB-IR: a Multi-modal Feature Pool module (MFP) and a Supplementary Feature Injector module (SFI). In the MFP module, the multi-receptive field convolution and the feature pyramid are deployed to capture contextual multi-scale features from RGB and IR images. These features are fused at different scales to complement foundation models for different RGB-IR semantic tasks. For the SFI module, a cross-attention mechanism is used to dynamically inject the required features into the pre-trained foundation model, which facilitates our UniRGB-IR framework with a robust feature representation ability. To inherit prior knowledge of the foundation model pre-trained on large-scale datasets, we utilize the adapter tuning paradigm instead of the full fine-tuning manner. We freeze the entire pre-trained weights and only optimize our proposed adapter. Consequently, UniRGB-IR can serve as a unified framework to achieve effective fine-tuning for various RGB-IR semantic tasks.
@@
  • We explore an efficient and universal framework called UniRGB-IR for RGB-IR semantic tasks. To the best of our knowledge, this is the first attempt to construct a unified framework for RGB-IR downstream tasks.
@@
  • We design a Multi-modal Feature Pool module and a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter dynamically injects the required features into the pre-trained model. These two modules can be efficiently fine-tuned with adapter tuning paradigm to complement the pre-trained foundation model with richer RGB-IR features for RGB-IR semantic tasks.
@@
  • We incorporate the vision transformer foundation model into the UniRGB-IR framework to evaluate the effectiveness of our method on RGB-IR downstream tasks, including RGB-IR object detection, RGB-IR semantic segmentation, and RGB-IR salient object detection. Extensive experimental results demonstrate that our methods can efficiently achieve superior performance on these RGB-IR downstream tasks.
The remainder of the article is organized as follows. In the Section II, we review the scientific literature related to our proposed approach. Section III introduces our proposed UniRGB-IR for RGB-Infrared semantic tasks. Experimental results are presented in Section IV, along with the comparison with state-of-the-art methods. Finally, we conclude with a discussion of our contribution in Section VI.
II Related Work
II-A Vision Foundation Models
@@
Recently, thanks to the powerful long-distance modeling capability, vision transformers (ViT) [7] have been widely used as the foundation model in many vision tasks to achieve competitive results. The original ViT is a plain, non-hierarchical structure for image classification. Based on it, ViTDet [20] also constructs a non-hierarchical model by incorporating the feature pyramid. However, the non-hierarchical structure lacks rich feature representation, resulting in unsatisfactory performance. Subsequently, various hierarchical transformers [21, 22, 23, 8] have been proposed for different downstream vision tasks. Swin Transformer [21] and Multiscale Vision Transformer [22] are designed based on the ViT model to explore multi-scale features to improve the performance of image classification and object detection tasks. Besides, PVT [24] performs global attention on the downsampled key and value maps for dense prediction. In this paper, we also introduce the ViT model as the pre-trained foundation model to construct a unified framework for RGB-IR semantic tasks.
II-B RGB-IR Semantic Tasks
@@
RGB-IR Object Detection. Zhang et al. [25] explore a two-stream SSD [26] structure to capture the contextual enhanced features for RGB-IR object detection. Besides, AR-CNN [27] is presented based on faster R-CNN [28] to align RGB and IR features. With the emergence of transformers, Yuan et al. [29] propose a complementary fusion transformer (CFT) module to achieve advanced detection results. Furthermore, C2Former [5] is a novel transformer block and can be incorporated into existing pre-trained models to increase intra- and inter-modality feature representations.
@@
RGB-IR Semantic Segmentation. MFNet [18] is proposed to incorporate infrared features into the RGB-based framework to perform RGB-IR semantic segmentation. Based on the transformer structure, Wu et al. [30] propose a novel CCFFNet to excavate complementary modality features for RGB-IR semantic segmentation. Moreover, CMX [31] is designed as a universal cross-modal fusion framework for RGB-IR semantic segmentation in an interactive fusion manner.
Figure 2: The overall architecture of our UniRGB-IR. In our framework, a ViT model with different numbers of ViT blocks is deployed as a foundation model, which is divided into $N$ (usually $N=4$) stages for feature interaction. Besides, a Multi-modal Feature Pool (MFP) is designed to extract contextual multi-scale RGB-IR features and a Supplementary Feature Injector (SFI) is explored to introduce these features into the ViT foundation model. During training, we freeze the entire ViT model weights and only optimize the MFP and SFI modules.
@@
RGB-IR Salient Object Detection. SwinNet [17] is designed based on the Swin Transformer to extract hierarchical information of each modality, which achieves impressive results. CAVER [32] introduces the transformer to rethink the bi-modal salient object detection from a sequence-to-sequence perspective, which increases the model interpretability. Recently, Zhou et al. [33] transfer a large amount of knowledge learned in the transformer-based network to lightweight WaveNet through the distillation method.
@@
The above methods attempt to design task-oriented structures to improve performance on corresponding downstream tasks. They either train the designed model from scratch or adopt a full fine-tuning strategy on a pre-trained model. Unlike these methods, we take a further step and deploy an adapter based on the pre-trained foundation model to unify the RGB-IR semantic tasks.
II-C Adapter
In the NLP field, Adapter [13] first fixes the original foundation backbone and introduces new modules into the transformer for task-specific fine-tuning, thereby effectively adapting the pre-trained backbone to downstream NLP tasks. Afterwards, Adapter has been widely studied in computer vision. ViT-Adapter [34] and Low-Rank Adapter [35] are designed to introduce a modest number of trainable parameters into the ViT and fine-tuning efficiently for dense prediction tasks. In addition, PC-Adapter [36] explores an attention-based adapter to preserve global shape knowledge for domain adaptation on point cloud data. Recently, Adapter is also used as a parameter-efficient training technology for vision-and-language tasks [37, 38].
In this paper, we aim to explore an adapter capable of converting the IR features into the RGB features to fit the pre-trained foundation model, which remains challenging to design this unified framework. Our UniRGB-IR is the first to propose utilizing adapter to unify RGB-IR downstream tasks.
III Method
III-A Overall Architecture
The overall framework of UniRGB-IR is illustrated in Figure 2, which consists of three parts: vision transformer model, Multi-modal Feature Pool (MFP) module, and Supplementary Feature Injector (SFI) module. In our framework, the ViT model is utilized as the pre-trained foundation model and frozen during the training process. Specifically, for the ViT model, the RGB image is directly fed into the patch embedding process to obtain the $D$-dimensional feature tokens, which are usually 1/16 of the original image resolution. To complement the richer features required for various RGB-IR downstream tasks, we feed the RGB and IR images into the MFP module to extract contextual multi-scale features from two modalities (eg. 1/8, 1/16 and 1/32 of the original image resolution). Afterwards, these richer features are dynamically injected into the features of ViT model through the SFI module, which can adaptively introduce the required RGB-IR features into the ViT model. To fully integrate the extracted features into the ViT model, we add the SFI module at the beginning of each stage. Thus, after $N$ stages of feature injection, the final features from ViT model can be leveraged for various RGB-IR downstream tasks.
@@ III-A Overall Architecture
Figure 3: Structure of the multi-modal feature pool (MFP) module. We explore the multiple perceptions to expand the receptive field of contextual feature extraction and utilize the feature pyramid to obtain the multi-scale features. In this module, Squeeze-Excitation(SE) attention is utilized to fuse two modality features. Thus, the contextual multi-scale features can be provided for the various RGB-IR downstream tasks.
III-B Multi-modal Feature Pool
To complement the rich feature representations for different RGB-IR downstream tasks, we introduce a multi-modal feature pool (MFP) module, including multiple perception and feature pyramid. The former can extract the contextual features with different convolution kernels, which achieve the long-distance modeling capability of CNNs. Different from existing works [39, 40] that increase the width or depth of the model, we efficiently achieve multi-receptive field perception in the channel dimension. As for feature pyramid, it can obtain multi-scale features to enhance the small object features. Thus, these two operations are connected in series to construct MFP module, as shown in Figure 3.
Specifically, for the input RGB ($H\times W\times 3$) and IR ($H\times W$) images, we first employ a stem block borrowed from ResNet to extract two modality features $F^{rgb}_{1}$ and $F^{ir}_{1}\in\mathbb{R}^{H/4\times W/4\times C}$. Then, these two features are split into four equal parts by utilizing the channel splitting technique [41]. To extract multi-receptive field perception, each part is subjected to convolution operations with different kernel sizes ($3\times 3$, $3\times 3$, $5\times 5$ and $7\times 7$). Then, we fuse each processed feature from two modalities by using a SE attention module (shown in Figure 3). Therefore, we concatenate each fused part to obtain RGB-IR contextual features $F_{fus}$. The above process can be represented as:
$F_{fus}=\varGamma_{k=1}^{4}(Fusion(W_{k}^{rgb}*f_{k}^{rgb},W_{k}^{ir}*f_{k}^{ir})),$ (1)
where $F_{fus}\in\mathbb{R}^{H/4\times W/4\times C}$, $f_{k}^{rgb}$ and $f_{k}^{ir}$ are the $k$-th part of $F_{1}^{rgb}$ and $F_{1}^{ir}$ features respectively, $W_{k}$ is the convolution with $k$-th kernel size, $\varGamma$ is the concatenation operation.
For the feature pyramid, a stack of three $3\times 3$ convolutions with $stride=2$ is conducted to downsample the size of the feature maps. Then, features of each scale are fed into a $1\times 1$ convolution to project the feature maps to $D$ dimensions. Therefore, we can obtain a set of multi-scale features {$F_{2}$, $F_{3}$, $F_{4}$} with 1/8, 1/16, and 1/32 of the original image resolution, respectively. Finally, we flatten and concatenate these features into feature tokens $F_{mfp}\in\mathbb{R}^{(\frac{HW}{8^{2}}+\frac{HW}{16^{2}}+\frac{HW}{32^{2}})\times D}$, which will be used as supplementary features for the ViT model.

### v4

# UniRGB-IR: A Unified Framework for Visible-Infrared Semantic Tasks via Adapter Tuning

@@ Abstract.
Semantic analysis on visible (RGB) and infrared (IR) images has gained significant attention due to their enhanced accuracy and robustness under challenging conditions including low-illumination and adverse weather. However, due to the lack of pre-trained foundation models on the large-scale infrared image datasets, existing methods prefer to design task-specific frameworks and directly fine-tune them with pre-trained foundation models on their RGB-IR semantic relevance datasets, which results in poor scalability and limited generalization. To address these limitations, we propose UniRGB-IR, a scalable and efficient framework for RGB-IR semantic tasks that introduces a novel adapter mechanism to effectively incorporate rich multi-modal features into pre-trained RGB-based foundation models. Our framework comprises three key components: a vision transformer (ViT) foundation model, a Multi-modal Feature Pool (MFP) module, and a Supplementary Feature Injector (SFI) module. The MFP and SFI modules cooperate with each other as an adpater to effectively complement the ViT features with the contextual multi-scale features. During training process, we freeze the entire foundation model to inherit prior knowledge and only optimize the MFP and SFI modules. Furthermore, to verify the effectiveness of our framework, we utilize the ViT-Base as the pre-trained foundation model to perform extensive experiments. Experimental results on various RGB-IR semantic tasks demonstrate that our method can achieve state-of-the-art performance. The codes are available at https://github.com/PoTsui99/UniRGB-IR.
RGB-IR semantic tasks; multi-modal fusion; adapters
Corresponding Author.
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10doi: 10.1145/3746027.3754806ccs: Computing methodologies Object detectionccs: Computing methodologies Image segmentationccs: Computing methodologies Scene understanding
1. Introduction
@@
Image semantic analysis in the visible images is a common practice in computer vision and has been widely used in a variety of vision tasks such as object detection (Yan et al., 2022; Bo et al., 2021), instance segmentation (Khan et al., 2020; Zhu et al., 2024), and semantic segmentation (Wang et al., 2025). However, visible cameras have demonstrated significant limitations in providing reliable imaging (Xu et al., 2017; Yuan and Wei, 2024), due to their restricted spectral bandwidth. This constraint becomes particularly problematic under low-illumination conditions and adverse weather scenarios. Therefore, infrared (IR) imaging, with its superior low-light adaptability, has been increasingly employed as complementary information to enhance visible modality performance (Xu et al., 2017; Zhao et al., 2024b). Afterwards, the joint use of RGB and IR images has been applied in more and more semantic analysis tasks.
@@
With the rapid advancements in physical computing capabilities, more and more general-purpose foundation backbones (Dosovitskiy et al., 2020; Xia et al., 2024) pre-trained on large-scale RGB-based datasets (e.g., ImageNet (Deng et al., 2009) and COCO (Lin et al., 2014)) have been designed for various tasks. Since these pre-trained models implicitly encode substantial prior knowledge, it is believed that pre-trained models can promote the improvement of downstream task performance and speed up training convergence. Therefore, fine-tuning the pre-trained foundation models (Wei et al., 2023; Shi et al., 2023) on their semantic relevance datasets has become a common paradigm to improve the performance on downstream tasks.
@@
However, due to the lack of pre-trained foundation models on the large-scale infrared image datasets, for RGB-Infrared semantic tasks, a straightforward and prevailing way (Zhang et al., 2021b; Liu et al., 2021a) is to use pre-trained RGB-based models and fine-tune them on their RGB-IR semantic relevance datasets as shown in Figure 1 (a). For example, RGB-IR object detectors (Zhang et al., 2021b; Yuan et al., 2022) utilize RGB-based models as strong baselines and fine-tune them to extract RGB and IR features. SwinNet (Liu et al., 2021c) employs transformer architecture to extract hierarchical features across RGB and infrared modalities, effectively detecting salient RGB-IR objects. Similarly, Ha et al. (Ha et al., 2017) establish a novel baseline by incorporating infrared modality features into an RGB-based framework and fine-tune it on the RGB-IR semantic segmentation benchmark. Among the efforts to achieve competitive results, these approaches still suffer from two major issues:
Limited model versatility and redundant parameter storage. Due to the lack of a unified framework for RGB-IR semantic tasks, different semantic tasks often require customized model structures tailored to specific objectives. However, task-specific customization often leads to poor model versatility, as each task demands its own dedicated model structure. Besides, maintaining multiple specialized models for different downstream tasks results in excessive parameter storage requirements.
Impairment of prior knowledge encoded in foundation models. Pre-trained foundation models are typically initialized using large-scale datasets, where they learn rich feature representations that capture general visual patterns. However, the full fine-tuning strategy indiscriminately updates the model parameters to adapt to the task-specific RGB-IR datasets, which often overrides the prior knowledge encoded in the pre-trained model. This will reduce the generalization potential of the fine-tuned model.
@@
Drawing inspiration from recent advances in adapters (Houlsby et al., 2019; Stickland and Murray, 2019), which are initially used to yield an extensible model to effectively exploit the potential representation of foundation models in the natural language processing (NLP) field, we develop an adapter to dynamically introduce extensive RGB-IR features into the pre-trained RGB-based foundation model. In our adapter, a feature extractor is designed to obtain rich RGB-IR features and a feature injector is proposed to adaptively introduce the required features for the foundation model. Without altering the original RGB-based foundation model, the robust pre-trained weights can be directly preserved to expedite training convergence and enable efficient fine-tuning for downstream tasks. Consequently, in this paper, we establish a Unified framework for RGB-IR semantic tasks, termed UniRGB-IR, as illustrated in Figure 1 (b).
Figure 1. Existing full fine-tuning methods vs. our UniRGB-IR framework. (a) Existing methods use pre-trained RGB-based foundation models and fully fine-tune them on their RGB-IR semantic relevance datasets. (b) We utilize the Adapter (Houlsby et al., 2019) to propose a unified framework, which can efficiently introduce richer RGB-IR features into the pre-trained foundation model for various semantic tasks.
Figure 2. The overall architecture of our UniRGB-IR. In our framework, a ViT model with different numbers of ViT blocks is deployed as a foundation model, which is divided into NN (usually $N=4$) stages for feature interaction. During training, we freeze the entire ViT model weights and only optimize the MFP and SFI modules.
Specifically, we design two core modules in our UniRGB-IR: a Multi-modal Feature Pool module (MFP) and a Supplementary Feature Injector module (SFI). For the MFP module, the multi-receptive field convolution and the feature pyramid are deployed to capture contextual multi-scale features from RGB and IR images. These features are fused at different scales to complement foundation models for different RGB-IR semantic tasks. As for the SFI module, we implement a cross-attention mechanism that dynamically injects essential features into the pre-trained foundation model, endowing our UniRGB-IR framework with robust feature representation capabilities. To inherit prior knowledge of the foundation model pre-trained on the large-scale datasets, we utilize the adapter tuning paradigm instead of the full fine-tuning manner. We freeze the pre-trained weights and only optimize the adapter. Consequently, our UniRGB-IR can serve as a unified framework to achieve effective fine-tuning for various RGB-IR semantic tasks.
@@
  • We explore an scalable and efficient framework called UniRGB-IR for RGB-IR semantic tasks. To the best of our knowledge, this is the first attempt to construct a unified framework for various RGB-IR downstream tasks.
@@
  • We design a Multi-modal Feature Pool module alongside a Supplementary Feature Injector module. The former extracts contextual multi-scale features from two modality images, and the latter dynamically injects the required features into the pre-trained model. These two modules can be efficiently fine-tuned with adapter tuning paradigm to complement the pre-trained foundation model with richer RGB-IR features for specific semantic task.
@@
  • We incorporate the vision transformer foundation model into the UniRGB-IR framework to evaluate the effectiveness of our method on RGB-IR semantic tasks, including RGB-IR object detection, RGB-IR semantic segmentation, and RGB-IR salient object detection. Extensive experimental results demonstrate that our methods can efficiently achieve superior performance on these downstream tasks.
2. Related Work
2.1. Vision Foundation Models
@@
Recently, thanks to the powerful long-distance modeling capability, vision transformers (ViT) (Dosovitskiy et al., 2020) have been widely used as the foundation model in many vision tasks to achieve competitive results. The original ViT is a plain, non-hierarchical structure for image classification. Based on it, ViTDet (Li et al., 2022a) also constructs a non-hierarchical model by incorporating the feature pyramid. However, the non-hierarchical structure lacks rich feature representation, resulting in unsatisfactory performance. Subsequently, various hierarchical transformers (Liu et al., 2021b; Fan et al., 2021; Li et al., 2023b; Xia et al., 2024) have been proposed for different downstream vision tasks. Swin Transformer (Liu et al., 2021b) and Multiscale Vision Transformer (Fan et al., 2021) are designed based on the ViT model to explore multi-scale features to improve the performance of image classification and object detection tasks. Besides, PVT (Wang et al., 2021) performs global attention on the downsampled key and value maps for dense prediction. In our work, we leverage the ViT model as our pre-trained foundation to build a unified framework for RGB-IR semantic tasks.
2.2. RGB-IR Semantic Tasks
@@
RGB-IR Object Detection. Zhang et al. (Zhang et al., 2019b) explore a two-stream SSD (Liu et al., 2016a) structure to capture the contextual enhanced features for RGB-IR object detection. Besides, AR-CNN (Zhang et al., 2019c) is presented based on Faster R-CNN (Ren et al., 2015) to align RGB and IR features. With the emergence of transformers, Yuan et al. (Yuan et al., 2024) propose a complementary fusion transformer (CFT) module to achieve advanced detection results. Furthermore, C2Former (Yuan and Wei, 2024) is a novel transformer block and can be incorporated into exited pre-trained models to increase intra- and inter-modality feature representations.
@@
RGB-IR Semantic Segmentation. MFNet (Ha et al., 2017) is proposed to incorporate infrared features into the RGB-based framework to perform RGB-IR semantic segmentation. Based on the transformer structure, Wu et al. (Wu et al., 2022a) propose a novel CCFFNet to excavate discriminative and complementary modality features for RGB-IR semantic segmentation. Moreover, CMX (Zhang et al., 2023a) is designed as a universal cross-modal fusion framework for RGB-IR semantic segmentation in an interactive fusion manner.
@@
RGB-IR Salient Object Detection. SwinNet (Liu et al., 2021c) is designed based on the Swin Transformer to extract hierarchical information of each modality, which achieves impressive results. CAVER (Pang et al., 2023) introduces the transformer to rethink the bi-modal salient object detection from a sequence-to-sequence perspective, which increases the model interpretability. Recently, Zhou et al. (Zhou et al., 2023a) transfer a large amount of knowledge learned in the transformer-based network to lightweight WaveNet through the distillation method.
@@
The above methods attempt to design task-oriented structures to improve performance on corresponding downstream tasks. They either train the designed model from scratch or adopt a full fine-tuning strategy on a pre-trained model. Unlike these methods, we deploy an adapter based on the pre-trained foundation model for various RGB-IR semantic tasks.
@@
Figure 3. Structure of the multi-modal feature pool (MFP) module. We explore multiple perceptions to expand the receptive field of contextual feature extraction and utilize the feature pyramid to obtain the multi-scale features.
2.3. Adapters
In the NLP field, Adapter (Houlsby et al., 2019) first fixes the original foundation backbone and introduces new modules into the transformer for task-specific fine-tuning, thereby effectively adapting the pre-trained backbone to downstream NLP tasks. Afterwards, Adapter has been widely studied in computer vision. ViT-Adapter (Chen et al., 2022a), Low-Rank Adapter (Yin et al., 2023) and Mona (Yin et al., 2025) are designed to introduce a modest number of trainable parameters into the ViT and fine-tuning efficiently for dense prediction tasks. In addition, PC-Adapter (Park et al., 2023) explores an attention-based adapter to preserve global shape knowledge for domain adaptation on point cloud data. Recently, Adapter is also used as a parameter-efficient training technology for vision-and-language tasks (Sung et al., 2022; Upadhyay et al., 2023).
In this paper, we aim to explore an adapter capable of converting the IR features into the RGB features to fit the pre-trained foundation model, which remains a challenge to design this unified framework. Our UniRGB-IR is the first to propose utilizing adapter for RGB-IR semantic tasks.
3. Method
3.1. Overall Architecture
The overall framework of UniRGB-IR is illustrated in Figure 2, which consists of three parts: vision transformer model, Multi-modal Feature Pool (MFP) module, and Supplementary Feature Injector (SFI) module. In our framework, the ViT model is utilized as the pre-trained foundation model and frozen during the training process. Specifically, for the ViT model, the RGB image is directly fed into the patch embedding process to obtain the $D$-dimensional feature tokens, which are usually 1/16 of the original image resolution. To complement the richer features required for various RGB-IR semantic tasks, we feed the RGB and IR images into the MFP module to extract contextual multi-scale features from two modalities (eg. 1/8, 1/16 and 1/32 of the original image resolution). Afterwards, these richer features are dynamically injected into the features of ViT model through the SFI module, which can adaptively introduce the required RGB-IR features into the ViT model. To fully integrate the extracted features into the ViT model, we add the SFI module at the beginning of each stage. Consequently, after $N$ stages of feature injection, the final features from ViT model can be leveraged for various RGB-IR semantic tasks.